### v1

# The Regular property of Invariant Rings over a Dedekind Domain

Abstract.
Let AA be a Dedekind domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing AA). Assume that $|G|$ is invertible in AA. We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Key words and phrases:
invariant rings, ring of integers of global fields, regular rings, group cohomology.
2020 Mathematics Subject Classification:
13A50, 13H05
1. Introduction
Consider the hypothesis below.
Hypothesis $1.$
Let AA be a Dedekind domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing AA). Assume that $|G|$ is invertible in AA.
In this article we establish the following result which is a variant of Theorem 7.2.1 in [1] for polynomial ring over a Dedekind domain.
Theorem 1.2.
Assume Hypothesis $1.$ Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Example 1.3.
The conclusion of Theorem 1.2 is true in the following cases.
  1. (1)
    AA is the ring of integers of a global function field i.e. a finite extension of $\mathbb{F}_{q}$(t) where $q$ is some power of a prime $p$, and $p\nmid|G|$.
  2. (2)
    AA is localization at a prime of the ring of integers of a number field and $|G|$ is non-zero in the residue field.
Remark 1.3.1.
We might wonder whether Theorem 1.2 holds if AA is a regular domain of higher dimension. However our technique fails when $\dim A\geq 2$, see 2.3.1.
2. Auxiliary Results
Consider the hypothesis below.
Hypothesis 2.1.
Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is a field). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$ with $|G|=r$ invertible in $\mathcal{O}/\pi\mathcal{O}$.
We can easily generalize Proposition 3.5 in [2] (which was for $n=2$) for any $n\in\mathbb{N}$ for our case. We provide the proof below for the sake of completeness.
Proposition $2.$
Assume Hypothesis 2.1. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is injective.
Proof.
We consider elements of $GL_{n}(\mathcal{O})$ and $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\pi\mathcal{O}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is rIrI which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $\mathcal{O}^{n}$ is noetherian module over $\mathcal{O}$, the map is an isomorphism. Thus $U\in GL_{n}(\mathcal{O})$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
The following is an integral lemma for proving Theorem 1.2.
Lemma 2.3.
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$ and any psuedo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
Proof.
We prove the result by induction on $n\geq 2$. Let $n=2$ and let $\{v_{1},v_{2}\}$ be a basis of $\mathcal{O_{\pi}}^{2}$ such that $\sigma(v_{1})=v_{1}$ and $\sigma(v_{2})=\lambda v_{2}$. We can assume $v_{1},v_{2}\in\mathcal{O}^{2}\backslash\pi\mathcal{O}^{2}$. Note $\lambda\neq 1$. Let $w_{1}=v_{1}$ and let $\{w_{1},w_{2}\}$ be a basis of $\mathcal{O}^{2}$. Let $v_{2}=\alpha w_{1}+\beta w_{2}$. Note that $\beta\neq 0$. Now $\sigma(v_{2})=\lambda v_{2}=\lambda\alpha w_{1}+\lambda\beta w_{2}$. Also we have $\sigma(v_{2})=\alpha\sigma(w_{1})+\beta\sigma(w_{2})=\alpha w_{1}+\beta\sigma(w_{2})$. Thus comparing both equations we have
(2.1) $\beta\sigma(w_{2})=(\lambda-1)\alpha w_{1}+\lambda\beta w_{2}$
Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$. Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid r$. As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$. By Proposition 2.2, the map $\lambda>\rightarrow(\mathcal{O}/\pi\mathcal{O})^{\times}$ is injective. As $\lambda\neq 1$, we have that the residue of $\lambda$ in $\mathcal{O}/\pi\mathcal{O}$ is $\bar{\lambda}\neq 1$. Thus $\lambda-1\not\in\pi\mathcal{O}$. So $\lambda-1\in\mathcal{O}^{\times}$.
We claim that $\beta\in\mathcal{O}^{\times}$. Assume on the contrary $\beta\in\pi\mathcal{O}$. Since $\alpha w_{1}+\beta w_{2}=v_{2}\not\in\pi\mathcal{O}^{2}$. Thus $\alpha\not\in\pi\mathcal{O}$. Going modulo $\pi$ in the Equation 2.1, we have $(\bar{\lambda}-1)\bar{\alpha}\bar{w_{1}}=0$ which is a contradiction. Thus we can write Equation 2.1 as $\sigma(w_{2})=\lambda w_{2}+\theta w_{1}$ for some $\theta\in\mathcal{O}$. Then letting $w^{\prime}_{2}=\frac{\theta}{\lambda-1}w_{1}+w_{2}$, we have that $\{w_{1},w^{\prime}_{2}\}$ is a basis and $\sigma(w^{\prime}_{2})=\lambda w^{\prime}_{2}$.
Let $n\geq 3$ and $\{v_{1},v_{2},\dots,v_{n}\}$ be a basis of $\mathcal{O}_{\pi}$ with $v_{i}\in\mathcal{O}^{n}\backslash\pi\mathcal{O}^{n}$ and $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$ and $\lambda\neq 1$. Let $E=\mathcal{O}v_{1}$ and $w_{1}=v_{1}$. We have the following exact sequence.
$0\rightarrow E\rightarrow W=\mathcal{O}^{n}\rightarrow\bar{W}=\mathcal{O}^{n-1}\rightarrow 0.$
Now $\sigma|_{\bar{W}}$ is also a psuedo-reflection. By induction we have a basis $\{\bar{w_{2}},\bar{w_{3}},\dots,\bar{w_{n}}\}$ of $\bar{W}$ with $\sigma(\bar{w_{i}})=\bar{w_{i}}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda\bar{w_{n}}$. Now $\{w_{1},w_{2},\dots,w_{n}\}$ is a basis of $W$ and $\sigma(w_{1})=w_{1}$ and $\sigma(w_{i})=w_{i}+a_{i}w_{1}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}+a_{n}w_{1}$. We will show that $a_{2}=a_{3}=\dots=a_{n-1}=0$. Consider $2\leq i\leq n-1$. Now $\sigma(w_{i})=w_{i}+a_{i}w_{1}$. Thus $\sigma^{j}(w_{i})=w_{i}+ja_{i}w_{1}$. Now $\sigma^{m}=id$ where $m=ord(\sigma)=ord(\lambda)$. Thus $w_{i}=w_{i}+ma_{i}w_{1}$. Hence $ma_{i}w_{1}=0$. As $m$ is invertible in $\mathcal{O}$, we have $a_{i}=0$. Also by Proposition 2.2, $\lambda-1$ is invertible in $\mathcal{O}$. Thus $w_{1},w_{2},\dots,w_{n-1},w^{\prime}_{n}$ is the required basis where $w^{\prime}_{n}=\frac{a_{n}}{\lambda-1}w_{1}+w_{n}$ and $\sigma(w^{\prime}_{n})=\lambda w^{\prime}_{n}$. ∎
Corollary 2.3.1.
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by psuedo-reflections.
Proof.
Follows from Proposition 2.2 and Lemma 2.3. ∎
Remark 2.3.1.
Our proof of Lemma 2.3 fails when $\mathcal{O}$ has dimension $\geq 2$. Suppose
$\mathcal{O}=k[X,Y]_{(X,Y)}$. Let $v=e_{1}/x^{2}+e_{2}/y^{2}\in K^{2}$ be such that $\sigma(v)=\lambda v$. Then note that for no choice of $t\in\mathcal{O}$ does $tv\in\mathcal{O}^{2}\setminus\mathfrak{m}\mathcal{O}^{2}$. So our induction fails at the base case $n=2$.
3. Proof of Theorem 1.2
Firstly we observe the following.
Hypothesis 3.1.
Consider the case in Hypothesis $1.$ Let $P$ be a maximal ideal of AA. Set $(\mathcal{O},(\pi))=(A_{P},PA_{P})$. Now $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}=K$. Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(\mathcal{O})$. Further since $|G|$ is invertible in AA, we have $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Thus for proving Theorem 1.2, it suffices to prove that $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Now let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$.
Now we give the proof of Theorem 1.2 by proving $S$ is regular in Hypothesis 3.1.
Proof.
Consider the short exact sequence
$0\rightarrow R\xrightarrow{\pi}R\rightarrow R/\pi R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{\pi}H^{1}(G,R)\rightarrow H^{1}(G,R/\pi R)\rightarrow\dots$
Let $K=ker(\pi)=Img(\delta)$ in the above sequence. Now we have that $K$ is $R^{G}/\pi R^{G}$-module. By Theorem 6.5.8 in [3] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. Hence $r\cdot K=0$. As $r$ is a unit in $R^{G}/\pi R^{G}$, we have $K=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\rightarrow 0.$
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.3.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections. Thus by Theorem 7.2.1 in [1], $(R/\pi R)^{G}$ is regular. Let $\mathfrak{m}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{m}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎