### v2

# The Regular property of Invariant Rings over Regular Domains

Abstract.
The main result of this paper is a generalization of the theorem of Chevalley-Shephard-Todd to the rings of invariants of pseudo-reflection groups over regular domains. More precisely, let $A$ be a regular domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$). Assume that $|G|$ is invertible in $A$. We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Key words and phrases:
invariant rings, regular domains, regular rings, group cohomology.
2020 Mathematics Subject Classification:
13A50, 13H05
1. Introduction
We have the famous Chevalley-Shephard-Todd Theorem (Theorem 7.2.1 in [1]) which states:
Theorem 1.1.
Let $K$ be a field. Consider a finite group $G\subseteq GL_{n}(K)$. Consider the ring
$K[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $K$). Assume that $|G|$ is invertible in $K$. We have $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
Mundelius had generalized Theorem 1.1 to the rings of invariants of pseudo-reflection groups over Dedekind domains (Theorem 3.8 in [2]).
In this paper we establish the following result which is a further generalization of Theorem 1.1 to the rings of invariants of pseudo-reflection groups over regular domains.
Theorem 1.2.
Let $A$ be a regular domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$. If $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections, then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
In Section 2, we prove Theorem 1.2 for the case when $A$ is a DVR as Theorem 2.2. This also leads to an alternate proof for the case of Dedekind domains considered by Mundelius in [2]. In Section 3, we prove Theorem 1.2 for the case when $A$ is a regular local ring as Theorem 3.2. Finally in Section 4, we complete the proof of Theorem 1.2.
2. Proof of Theorem 1.2 for DVRs
Consider the hypothesis below.
Hypothesis 2.1.
Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$. Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $\mathcal{O}$). Assume that $|G|$ is invertible in $\mathcal{O}$ i.e. $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
In this section, we will prove the following.
Theorem 2.2.
Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by psuedo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
We can easily generalize Proposition 3.5 in [3] (which was for $n=2$) for any $n\in\mathbb{N}$ for our case. We provide the proof below for the sake of completeness.
Proposition 2.3.
Assume Hypothesis 2.1. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is injective.
Proof.
We consider elements of $GL_{n}(\mathcal{O})$ and $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\pi\mathcal{O}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is $rI$ which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $\mathcal{O}^{n}$ is noetherian module over $\mathcal{O}$, the map is an isomorphism. Thus $U\in GL_{n}(\mathcal{O})$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
The following is an integral lemma for proving Theorem 2.2.
Lemma 2.4.
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$ and any psuedo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
Proof.
We prove the result by induction on $n\geq 2$. Let $n=2$ and let $\{v_{1},v_{2}\}$ be a basis of $\mathcal{O_{\pi}}^{2}$ such that $\sigma(v_{1})=v_{1}$ and $\sigma(v_{2})=\lambda v_{2}$. We can assume $v_{1},v_{2}\in\mathcal{O}^{2}\backslash\pi\mathcal{O}^{2}$. Note $\lambda\neq 1$. Let $w_{1}=v_{1}$ and let $\{w_{1},w_{2}\}$ be a basis of $\mathcal{O}^{2}$. Let $v_{2}=\alpha w_{1}+\beta w_{2}$. Note that $\beta\neq 0$. Now $\sigma(v_{2})=\lambda v_{2}=\lambda\alpha w_{1}+\lambda\beta w_{2}$. Also we have $\sigma(v_{2})=\alpha\sigma(w_{1})+\beta\sigma(w_{2})=\alpha w_{1}+\beta\sigma(w_{2})$. Thus comparing both equations we have
(2.1) $\beta\sigma(w_{2})=(\lambda-1)\alpha w_{1}+\lambda\beta w_{2}$
Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$. Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid r$. As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$. By Proposition 2.3, the map $\lambda>\rightarrow(\mathcal{O}/\pi\mathcal{O})^{\times}$ is injective. As $\lambda\neq 1$, we have that the residue of $\lambda$ in $\mathcal{O}/\pi\mathcal{O}$ is $\bar{\lambda}\neq 1$. Thus $\lambda-1\not\in\pi\mathcal{O}$. So $\lambda-1\in\mathcal{O}^{\times}$.
We claim that $\beta\in\mathcal{O}^{\times}$. Assume on the contrary $\beta\in\pi\mathcal{O}$. Since $\alpha w_{1}+\beta w_{2}=v_{2}\not\in\pi\mathcal{O}^{2}$. Thus $\alpha\not\in\pi\mathcal{O}$. Going modulo $\pi$ in the Equation 2.1, we have $(\bar{\lambda}-1)\bar{\alpha}\bar{w_{1}}=0$ which is a contradiction. Thus we can write Equation 2.1 as $\sigma(w_{2})=\lambda w_{2}+\theta w_{1}$ for some $\theta\in\mathcal{O}$. Then letting $w^{\prime}_{2}=\frac{\theta}{\lambda-1}w_{1}+w_{2}$, we have that $\{w_{1},w^{\prime}_{2}\}$ is a basis and $\sigma(w^{\prime}_{2})=\lambda w^{\prime}_{2}$.
Let $n\geq 3$ and $\{v_{1},v_{2},\dots,v_{n}\}$ be a basis of $\mathcal{O}_{\pi}$ with $v_{i}\in\mathcal{O}^{n}\backslash\pi\mathcal{O}^{n}$ and $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$ and $\lambda\neq 1$. Let $E=\mathcal{O}v_{1}$ and $w_{1}=v_{1}$. We have the following exact sequence.
$0\rightarrow E\rightarrow W=\mathcal{O}^{n}\rightarrow\bar{W}=\mathcal{O}^{n-1}\rightarrow 0.$
Now $\sigma|_{\bar{W}}$ is also a psuedo-reflection. By induction we have a basis $\{\bar{w_{2}},\bar{w_{3}},\dots,\bar{w_{n}}\}$ of $\bar{W}$ with $\sigma(\bar{w_{i}})=\bar{w_{i}}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda\bar{w_{n}}$. Now $\{w_{1},w_{2},\dots,w_{n}\}$ is a basis of $W$ and $\sigma(w_{1})=w_{1}$ and $\sigma(w_{i})=w_{i}+a_{i}w_{1}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}+a_{n}w_{1}$. We will show that $a_{2}=a_{3}=\dots=a_{n-1}=0$. Consider $2\leq i\leq n-1$. Now $\sigma(w_{i})=w_{i}+a_{i}w_{1}$. Thus $\sigma^{j}(w_{i})=w_{i}+ja_{i}w_{1}$. Now $\sigma^{m}=id$ where $m=ord(\sigma)=ord(\lambda)$. Thus $w_{i}=w_{i}+ma_{i}w_{1}$. Hence $ma_{i}w_{1}=0$. As $m$ is invertible in $\mathcal{O}$, we have $a_{i}=0$. Also by Proposition 2.3, $\lambda-1$ is invertible in $\mathcal{O}$. Thus $w_{1},w_{2},\dots,w_{n-1},w^{\prime}_{n}$ is the required basis where $w^{\prime}_{n}=\frac{a_{n}}{\lambda-1}w_{1}+w_{n}$ and $\sigma(w^{\prime}_{n})=\lambda w^{\prime}_{n}$. ∎
Corollary 2.4.1.
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by psuedo-reflections.
Proof.
Follows from Proposition 2.3 and Lemma 2.4. ∎
Now we give the proof of Theorem 2.2.
Proof.
Let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
$0\rightarrow R\xrightarrow{\pi}R\rightarrow R/\pi R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{\pi}H^{1}(G,R)\rightarrow H^{1}(G,R/\pi R)\rightarrow\dots$
By Theorem 6.5.8 in [4] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. As $r$ is a unit in $\mathcal{O}=(R^{G})_{0}$, we have that it is a unit in $R^{G}$. Thus $H^{1}(G,R)=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\rightarrow 0.$
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections. Thus by Theorem 1.1, $(R/\pi R)^{G}$ is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
3. Proof of Theorem 1.2 for regular local rings
Consider the hypothesis below.
Hypothesis 3.1.
Let $(A,\mathfrak{m})$ be a regular local ring and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$.
In this section, we will prove the following.
Theorem 3.2.
Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Now consider the following.
Hypothesis 3.3.
Consider the case in Hypothesis 3.1. Let $d=\dim(A)$. Now Theorem 3.2 for the case $d=0$ (i.e. $A$ is a field) follows from Theorem 1.1. Thus we can assume $d\geq 1$. Let $\mathfrak{m}=(x_{1},\dots,x_{d})$. Set $x=x_{1}$ and let $\mathfrak{p}$ be the prime ideal $(x)$.
We have an analogous result to Proposition 2.3 for regular local rings. Essentially the same proof goes through but we provide it for the sake of completeness.
Lemma 3.4.
Assume Hypothesis 3.3. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective.
Proof.
We consider elements of $GL_{n}(A)$ and $GL_{n}(A/\mathfrak{p})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\mathfrak{p}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(A/\mathfrak{p})$ is $rI$ which is invertible in $GL_{n}(A/\mathfrak{p})$ since $r$ is invertible in $A/\mathfrak{p}$. Consider the map $U:A^{n}\rightarrow A^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $A^{n}$ is noetherian module over $A$, the map is an isomorphism. Thus $U\in GL_{n}(A)$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
Now we give the proof of Theorem 3.2 with notations as in Hypothesis 3.3.
Proof.
We prove the result by induction on $d\geq 1$. Now the result for the case $d=1$ (i.e. $A$ is a DVR) follows from Theorem 2.2. So assume $d\geq 2$ and that the result is known for all regular local rings of dimension d1d-1.
As $A$ is a regular local ring we have that $A_{\mathfrak{p}}$ is a regular local ring of dimension $1$, i.e. a DVR which has $K$ as its field of fractions. Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$. Now $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections. Note that we have a commutative diagram of ring homomorphisms,
[…]
where $\mu$ is the standard ring homomorphism of $A/\mathfrak{p}$ to its field of fractions which is injective. As we have the above commutative diagram, we also have the below commutative diagram of group homomorphisms with $\mu$ being injective.
$GL_{n}(A)$$GL_{n}(A/\mathfrak{p})$$GL_{n}(A_{\mathfrak{p}})$$GL_{n}(\kappa(\mathfrak{p}))$$\scriptstyle{\iota}$$\scriptstyle{\gamma}$$\scriptstyle{\mu}$$\scriptstyle{\gamma^{\prime}}$
We have that $\gamma|_{G}:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective by Lemma 3.4. Thus we have that $\mu\circ\gamma|_{G}$ is injective. Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$. So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections. By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ to be regular. Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
$0\rightarrow R\xrightarrow{x}R\rightarrow R/\mathfrak{p}R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{x}H^{1}(G,R)\rightarrow H^{1}(G,R/\mathfrak{p}R)\rightarrow\dots$
By Theorem 6.5.8 in [4] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. As $r$ is a unit in $A=(R^{G})_{0}$, we have that it is a unit in $R^{G}$. Thus $H^{1}(G,R)=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\rightarrow 0.$
Thus $R^{G}/\mathfrak{p}R^{G}\cong(R/\mathfrak{p}R)^{G}$ which is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $x\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\mathfrak{p}R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
4. Proof of Theorem 1.2 for regular domains
Now we complete the proof of Theorem 1.2.
Proof.
Now $A$ is a regular domain and $K$ is its field of fractions. Let $P$ be a maximal ideal of $A$. So $(A_{P},PA_{P})$ is a regular local ring with $K$ as the field of fractions. Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(A_{P})$. Further since $|G|$ is invertible in $A$, we have $|G|$ is invertible in $A_{P}$. Thus for proving Theorem 1.2, it suffices to prove that $(A_{P}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular which follows from Theorem 3.2. ∎
Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius.