### v3

# The Regular property of Invariant Rings over Regular Domains

Abstract.
The main result of this paper is a generalization of the theorem of Chevalley-Shephard-Todd to the rings of invariants of pseudo-reflection groups over regular domains. More precisely, let $A$ be a regular domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$). Assume that $|G|$ is invertible in $A$. We prove that $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Key words and phrases:
invariant rings, regular domains, regular rings, group cohomology.
2020 Mathematics Subject Classification:
13A50, 13H05
1. Introduction
We begin by recalling the notion of pseudo-reflections.
Definition 1.1 (§7.1, [1]).
Let $K$ be a field. An element $\sigma\in GL_{n}(K)$ is called a pseudo-reflection if $ord(\sigma)$ is finite and $dim(ker(\sigma-1))=n-1$, i.e. whose fixed points have codimension one.
We have the famous Chevalley-Shephard-Todd Theorem (Theorem 7.2.1 in [1]) which states:
Theorem 1.2.
Let $K$ be a field. Consider a finite group $G\subseteq GL_{n}(K)$. Consider the ring
$K[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $K$). Assume that $|G|$ is invertible in $K$. We have $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
Mundelius had generalized Theorem 1.2 to the rings of invariants of pseudo-reflection groups over Dedekind domains (Theorem 3.8 in [3]).
In this paper we establish the following result which is a further generalization of Theorem 1.2 to the rings of invariants of pseudo-reflection groups over regular domains.
Theorem 1.3.
Let $A$ be a regular domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$. We have $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
One direction of the above theorem follows from Theorem 1.2. Here’s the proof.
Proof.
Suppose $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Now $G$ acts linearly on the ring $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$) and $|G|$ is invertible in $A$. As $K$ is the field of fractions of $A$, so $G$ also acts linearly on the ring $K[X_{1},X_{2},\dots,X_{n}]$ (fixing $K$) and $|G|$ is invertible in $K$. Now as $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a localization of $(A[X_{1},X_{2},\dots,X_{n}])^{G}$, we have that $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Therefore by Theorem 5.3.3 in [4] and by Theorem 19.2 in [2], $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over the field $K$. Hence by Theorem 1.2, $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections . ∎
The subsequent sections of the paper are dedicated to proving the other direction. In Section 2, we prove Theorem 1.3 for the case when $A$ is a DVR as Theorem 2.2. This also leads to an alternate proof for the case of Dedekind domains considered by Mundelius in [3]. In Section 3, we prove Theorem 1.3 for the case when $A$ is a regular local ring as Theorem 3.2. Finally in Section 4, we complete the proof of Theorem 1.3.
2. Proof of Theorem 1.3 for DVRs
Consider the hypothesis below.
Hypothesis 2.1.
Assume that $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}$ is the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$. Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and assume that $G$ acts linearly on the ring (fixing $\mathcal{O}$). Assume that $|G|$ is invertible in $\mathcal{O}$ i.e. $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
In this section, we will prove the following.
Theorem 2.2.
Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by pseudo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
We can easily generalize Proposition 3.5 in [5] (which was for $n=2$) for any $n\in\mathbb{N}$ for our case. We provide the proof below for the sake of completeness.
Proposition 2.3.
Assume Hypothesis 2.1. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is injective.
Proof.
We consider elements of $GL_{n}(\mathcal{O})$ and $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\pi\mathcal{O}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is $rI$ which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $\mathcal{O}^{n}$ is noetherian module over $\mathcal{O}$, the map is an isomorphism. Thus $U\in GL_{n}(\mathcal{O})$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
The following is an integral lemma for proving Theorem 2.2.
Lemma 2.4.
Assume Hypothesis 2.1. Suppose that $G$ is generated by pseudo-reflections in
$GL_{n}(\mathcal{O}_{\pi})$. Then for $n\geq 2$ and any pseudo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
Proof.
Let $\{v_{1},\ldots v_{n}\}$ be a basis of $(\mathcal{O}_{\pi})^{n}$ such that $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$.
Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$. Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid|G|$. As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$. By Proposition 2.3, the map $\lambda>\rightarrow(\mathcal{O}/\pi\mathcal{O})^{\times}$ is injective. As $\lambda\neq 1$, we have that the residue of $\lambda$ in $\mathcal{O}/\pi\mathcal{O}$ is $\bar{\lambda}\neq 1$. Thus $\lambda-1\not\in\pi\mathcal{O}$. So $\lambda-1\in\mathcal{O}^{\times}$.
Consider the endomorphism $\epsilon=(\lambda-1)^{-1}(\sigma-id)\colon\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By considering $\epsilon_{\pi}$ it follows that $\epsilon$ is an idempotent. So $\mathcal{O}^{n}=\ker\epsilon\oplus\ker(1-\epsilon)$. We note that $\operatorname{rank}(\ker(\epsilon))=n-1$ and $\operatorname{rank}(\ker(1-\epsilon))=1$. On $\ker(\epsilon)$ the map $\sigma$ acts as identity while on $\ker(1-\epsilon)$ the map $\sigma$ acts as multiplication by $\lambda$. The result follows. ∎
Corollary 2.4.1.
Assume Hypothesis 2.1. Suppose that $G$ is generated by pseudo-reflections in $GL_{n}(\mathcal{O}_{\pi})$. Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by pseudo-reflections.
Proof.
Follows from Proposition 2.3 and Lemma 2.4. ∎
Now we give the proof of Theorem 2.2.
Proof.
Let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
$0\rightarrow R\xrightarrow{\pi}R\rightarrow R/\pi R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{\pi}H^{1}(G,R)\rightarrow H^{1}(G,R/\pi R)\rightarrow\dots$
By Theorem 6.5.8 in [6] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. As $r$ is a unit in $\mathcal{O}=(R^{G})_{0}$, we have that it is a unit in $R^{G}$. Thus $H^{1}(G,R)=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\rightarrow 0.$
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by pseudo-reflections. Thus by Theorem 1.2, $(R/\pi R)^{G}$ is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
3. Proof of Theorem 1.3 for regular local rings
Consider the hypothesis below.
Hypothesis 3.1.
Assume that $(A,\mathfrak{m})$ is a regular local ring and $K$ is its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and assume that $G$ acts linearly on the ring (fixing $A$). Assume that $|G|$ is invertible in $A$.
In this section, we will prove the following.
Theorem 3.2.
Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Now consider the following.
Hypothesis 3.3.
Consider the assumptions in Hypothesis 3.1. Let $d=\dim(A)$. Now Theorem 3.2 for the case $d=0$ (i.e. $A$ is a field) follows from Theorem 1.2. Thus we can assume $d\geq 1$. Let $\mathfrak{m}=(x_{1},\dots,x_{d})$. Set $x=x_{1}$ and let $\mathfrak{p}$ be the prime ideal $(x)$.
We have an analogous result to Proposition 2.3 for regular local rings. Essentially the same proof goes through but we provide it for the sake of completeness.
Lemma 3.4.
Assume Hypothesis 3.3. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective.
Proof.
We consider elements of $GL_{n}(A)$ and $GL_{n}(A/\mathfrak{p})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\mathfrak{p}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(A/\mathfrak{p})$ is $rI$ which is invertible in $GL_{n}(A/\mathfrak{p})$ since $r$ is invertible in $A/\mathfrak{p}$. Consider the map $U:A^{n}\rightarrow A^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $A^{n}$ is noetherian module over $A$, the map is an isomorphism. Thus $U\in GL_{n}(A)$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
Now we give the proof of Theorem 3.2 with notations as in Hypothesis 3.3.
Proof.
We prove the result by induction on $d\geq 1$. Now the result for the case $d=1$ (i.e. $A$ is a DVR) follows from Theorem 2.2. So assume $d\geq 2$ and that the result is known for all regular local rings of dimension d1d-1.
As $A$ is a regular local ring we have that $A_{\mathfrak{p}}$ is a regular local ring of dimension $1$, i.e. a DVR which has $K$ as its field of fractions. Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$. Now $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections. Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by pseudo-reflections. Note that we have a commutative diagram of ring homomorphisms,
[…]
where $\mu$ is the standard ring homomorphism of $A/\mathfrak{p}$ to its field of fractions which is injective. As we have the above commutative diagram, we also have the below commutative diagram of group homomorphisms with $\mu$ being injective.
$GL_{n}(A)$$GL_{n}(A/\mathfrak{p})$$GL_{n}(A_{\mathfrak{p}})$$GL_{n}(\kappa(\mathfrak{p}))$$\scriptstyle{\iota}$$\scriptstyle{\gamma}$$\scriptstyle{\mu}$$\scriptstyle{\gamma^{\prime}}$
We have that $\gamma|_{G}:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective by Lemma 3.4. Thus we have that $\mu\circ\gamma|_{G}$ is injective. Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$. So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by pseudo-reflections. By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ is regular. Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
$0\rightarrow R\xrightarrow{x}R\rightarrow R/\mathfrak{p}R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{x}H^{1}(G,R)\rightarrow H^{1}(G,R/\mathfrak{p}R)\rightarrow\dots$
By Theorem 6.5.8 in [6] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. As $r$ is a unit in $A=(R^{G})_{0}$, we have that it is a unit in $R^{G}$. Thus $H^{1}(G,R)=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\rightarrow 0.$
Thus $R^{G}/\mathfrak{p}R^{G}\cong(R/\mathfrak{p}R)^{G}$ which is regular. Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$. Since $x\not\in\mathfrak{M}^{2}$. Thus $R^{G}/\mathfrak{p}R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
4. Proof of Theorem 1.3 for regular domains
Now we complete the proof of Theorem 1.3.
Proof.
Now $A$ is a regular domain and $K$ is its field of fractions. Let $P$ be a maximal ideal of $A$. So $(A_{P},PA_{P})$ is a regular local ring with $K$ as the field of fractions. Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(A_{P})$. Further since $|G|$ is invertible in $A$, we have $|G|$ is invertible in $A_{P}$. Thus for proving Theorem 1.3, it suffices to prove that $(A_{P}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular which follows from Theorem 3.2. ∎
5. Conflict of interest statement
The authors have no conflict of interest to declare that are relevant to this article.
6. Data availability statement
The manuscript has no associated data.
Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius. We also thank the two referees for many pertinent comments.